Database last updated on Sun, 01 Feb 2026 22:40:50 +0000 by localhost
Instrument constant changes for the week before 2023304 (3 sec, 0 rows)
HG burned out? (AP results when HG is ON) (0 sec, 17 rows)
Measurements and tests summary.
| _jday | _loc | _serial | _skc | aode | ds | ed | hg | sl | um | uv | zs |
| 304 | TORONTO | #015 | Next schedule will be to3mk2um (now to3mk2um ZS= 100 ZS= 100),to3mk2um,Waiting until ZA = 100 now 98.6 | 10 | 10 | 2 | 59 | 18 | 114 | 65 | 28 |
(0 sec)
Dead time (0 sec)
Nothing to report.
Standard lamp R6 (1 sec)
| File name | Count | Serial | Expected (+/-20) | Measured | 30-day history |
| B30423.015 | 18 | #015 | 1990 | 1803 |